OpenStudy (anonymous):

find the value of y for the given values of x & z. y=x+z x=6, z=17

OpenStudy (anonymous):

would this just be y=6+17

OpenStudy (anonymous):

y=23

OpenStudy (anonymous):

Yep, correct. Just replace x by a '6' and z by a '17' then y = x+z = 6+17 = 23

OpenStudy (anonymous):

ok thanks just thought i was looking to far into it

OpenStudy (anonymous):

no problem :)